OpenStudy (anonymous):

If f(x) = 6x + 3 and g(x) = 5x - 4, what is f(g(x))?

OpenStudy (anonymous):

f(g(x)) = f(5x - 4) = 6(5x - 4) + 3 = 30x - 24 + 3 = 30x - 21

OpenStudy (anonymous):

so the answer is 30x-21

OpenStudy (anonymous):

i worked it out and thats wht i got

OpenStudy (anonymous):

is it right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

thanks:D